OpenStudy (anonymous):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (anonymous):

@green_1

OpenStudy (anonymous):

@carlyleukhardt

OpenStudy (anonymous):

@Missiey

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

a

OpenStudy (anonymous):

what else

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

d

OpenStudy (theedwardsfamily):

What do you think @goopyfish908 ?

OpenStudy (anonymous):

A and B

OpenStudy (anonymous):

am i right, or is it a and d

OpenStudy (theedwardsfamily):

My explaination: 0 × 0 = 0 0 × 1 = 0 1 × 1 = 1 Which means multiplication is closed under {0, 1} 1 ÷ 1 = 1 0 ÷ 1 = 0 Division is not closed under {0, 1} 1 + 1 = 2 Addition is not closed under {0, 1} 0 - 1 = -1 Subtraction is not closed under {0, 1} either So it makes sense to only be A. Multiplication which is closed under {0, 1} Comments (0)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

@Missiey

OpenStudy (mathstudent55):

First, do you know what the set of whole numbers is?

OpenStudy (mathstudent55):

Then, choose examples that show addition of only whole numbers.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

@carlyleukhardt

OpenStudy (mathstudent55):

When a set of numbers is closed for an operation, it means that if you take any numbers form that set and do that operation, the answer is also a part of the set. Let's do an example. Question: Is the set {0, 1) closed for addition? There are two elements in this set, 0 and 1. What are all the different additions possible with these two numbers? They are: 0 + 0 0 + 1 1 + 0 1 + 1 What are the results of the four additions above? 0 + 0 = 0 0 + 1 = 1 1 + 0 = 1 1 + 1 = 2 The first three additions have results of 0 or 1, both of which are in the set {0, 1}, but the fourth addition, 1 +1 has a result of 2. Since 2 is not in the set {0, 1}, this means the set {0, 1} is not closed for addition because there is at least one addition using elements of the set that gives a result that is not an element of the set.

OpenStudy (mathstudent55):

Your problem uses only the set {0, 1}. Then is asks if the set {0, 1} is closed for multiplication, division, addition, and subtraction. I showed you above that the set is not closed for addition. Now do a similar thing for the other operations.